# Controlling quantum vortex dynamics and vortex-antivortex annihilation in Bose-Einstein condensates with optical lattices

### v1

@@ I Introduction and summary
A subject that apparently has received very little attention is the behavior of vortex dipoles. Such configurations have been experimentally obtained in BEC in the work of Refs.neely . However topics such as the interaction of such dipoles and the ensuing dynamics are still unexplored. Our study provide an insight into these issues, and how they are modified by optical lattice.
@@
Vortices in the SS state of matter in dipolar bosons have been theoretically addressed gallemi ; casotti ; gallemi1 ; Anc21 in recent years and some relevant phenomena have been uncovered, like a reduced angular momentum associated to the quantum of circulation gallemi . A peculiar behavior has been found Anc21 in the dynamic of a vortex dipole, i.e. a pair of vortices of opposite chirality, in a SS. In a uniform superfluid, when the distance $l$ between the two vortices is larger than the healing length $\xi$, the pair is a stable entity and moves with a constant velocity as a rigid body in a direction perpendicular to the vector $\vec{l}$ joining the two vortices. Contrary to this, in dipolar bosons in a SS state it was found that the two vortices moves by jumps between equilibrium sites in the lattice and approach each other until the two vortices annihilate themselves in a short time. At the basis of this behavior is a basic property of a vortex in a SS: it is energetically favorable to have its core at a discrete set of positions, the locus of minimum density. The natural question is if such properties are specific of a SS state or if they are generic ones when a density modulation is present, whatever is its origin. To answer this question we address in the present paper the study of bosons in an external periodic potential like that produced by light standing waves. We study the ground state and vortices of the standard model of BEC, point-like bosons with a contact interaction, in the case of 87Rb with the Gross-Pitaevskii equation in an external periodic potential $V_{ext}(\vec{r})$. We consider the case of modulation in one dimension (1D) and in two dimensions (2D) for three lattices: square, triangular and honeycomb.
@@
If the two vortices of the dipole are located in the same channel they do not translate but they move one against the other until they annihilate and the excitation energy goes into phonons of the superfluid. If the two vortices are located in adjacent channels they move along these channels and the motion is a composition of a uniform translation and of an oscillation. In all cases the translation velocity of the pair differs from that expected for a uniform superfluid and it can even invert the direction of motion for large amplitude of the optical potential. A combination of translation and oscillation is found also for a single vortex in a trap in presence of a 1D OL.. The frequency of this oscillation is the same as in the case of a vortex dipole moving along two neighboring channels. Therefore this oscillatory motion seems to be an intrinsic and novel character of the motion of a vortex moving along a channel. Notice that this oscillatory motion is quite different from that of a massive particle around a minimum energy position because the vortex oscillation is one-side, i.e. the vortex never crosses the position of the minimum energy and it remains on the side from which it started.
@@ II Method
where $V({\bf r})$ and $\rho({\bf r})=|\psi({\bf r})|^{2}$ represent the external potential and the boson number density, respectively. The coupling constant is $g=4\pi a_{s}\hbar^{2}/m$, $m$ being the atomic mass. The number density $\rho$ is normalized such that $\int_{V}\rho({\bf r})\,{\rm d}{\bf r}=N$ where $N$ is the total number of atoms. As model system we take 87Rb atoms and the scattering length $a_{s}$ describing the (repulsive) Rb-Rb interaction is $a_{s}=100.4\,a_{0}$ marte .
@@
Alternatively, one could estimate the superfluid fraction from leg70
@@
where $<n(x)>$ is the number density of the ground-state configuration averaged over the transverse y-z directions. We find that the values of $f_{s}$ from the two definitions agree with each other to within $1\%$.
@@ III.1 Single vortex properties
The flow field of a linear vortex has a long-range character, $\sim 1/r$, where $r$ is the distance from the position of the vortex core. We have imposed antiperiodic boundary conditions Pi07 in order that the condition of no flow across the boundary of the computational cell is satisfied Anc21 . This is equivalent to sum over the phases of an infinite array of vortex-antivortex, i.e. a vortex of opposite chirality is present in each nearest neighbor cell of the computation cell sadd . Equation (7) can be easily generalized to accommodate a vortex array made of an arbitrary number of vortices and/or antivortices Anc18 . The case of a pair of vortices with opposite circulation (vortex dipole) will be considered in in the following. In the case of a vortex dipole we impose however the usual periodic boundary conditions since the flow fields of the two vortices tend to cancel at the cell boundaries.
@@
The equilibrium vortex structure at one of such sites is shown in Fig.3 for the case $\tilde{V}=2$. The streamlines for the velocity field $\vec{v}=\hbar\nabla\phi/m$ ($\phi$ being the phase of the wavefunction) are shown in Fig.4. Notice the strong deformations with respect to the circular patterns expected from a vortex in a homogeneous system. In order to display these deformations more clearly in Fig.5 we also show the streamlines of the modified phase ${\tilde{\phi}}$ as defined in Eq.(8) further on. ${\tilde{\phi}}$ represents the deviation of the phase from the imprinted phase in Eq.(7).
@@ III.2 Vortex dipole properties
From the equation for the velocity of a vortex within the GP equation we can understand at a qualitative level some aspects of our results. The flow field of one vortex on the other is in the transverse direction of the channel, i.e. in the region where the density of the superfluid is increasing. This variation of density gives rise to a component of the velocity of the vortex in the longitudinal direction along the bottom of the channel. Following Ref.simula we write the vortex wave function $\psi_{v}(\vec{r},t)$ with the core located at position $\vec{r}_{0}$ at time $t$ as
@@
where $\tilde{\rho}(\vec{r},t)$ and $\tilde{\phi}(\vec{r},t)$ represent deviations of the modulus and of the phase from the ideal gas form, the first factor in eq.(8). $\tilde{\rho}(\vec{r},t)$ and $\tilde{\phi}(\vec{r},t)$ contain the following effects: the contribution due to the presence of other vortices, the effect of an external potential and, finally, the deviation of the $\psi_{v}$ from the ideal gas form due to the inter-atomic interactions. On the basis of the Gross-Pitaevskii equation here adopted the velocity of the vortex is simula
$\tilde{\vec{v}}\equiv\frac{\hbar}{m}(\nabla\tilde{\phi}-\hat{\kappa}\times\nabla log\tilde{\rho})$ (9)
@@ III.3 Single vortex in a trap
$V_{box}(x,y)=U_{0}(1-\frac{1}{e^{(R-R_{0})/\sigma}+1})$ (10)
@@
We perform the computation with the external potential $V(\vec{r})=V_{OL}(x,y)+V_{trap}(x,y)$ and $V_{OL}$ is such that the center of the trap is a maximum of the OL, i.e. the center of the trap is an equilibrium position for the vortex core. We find that the vortex remains immobile if initially it is located at or very near the center of the trap, in the low density channel passing through it. If the initial position of the vortex is off-center along the y direction by a larger amount, however, the vortex is found to migrate towards the edge of the trap and disappears once it approaches the trap edge.
@@ IV.3 Vortex dipole properties
The energy released immediately after the annihilation goes into excitations of the system. In order to gain some insights on the character of the excitations we computed the spectral density of the kinetic energy of the superfluid velocity field, decomposing it into compressible and incompressible parts tsubota ; spectra ; bradley . Briefly, one splits the density-weighted velocity field $\vec{u}(\vec{r},t)\equiv\sqrt{\rho(\vec{r},t)}\vec{v}(\vec{r},t)$ into a compressible (C) and an incompressible (I) part, $\vec{u}(\vec{r},t)=\vec{u}^{I}(\vec{r},t)+\vec{u}^{C}(\vec{r},t)$, such that $\vec{\nabla}\cdot\vec{u}^{I}(\vec{r},t)=0$ and $\vec{\nabla}\wedge u^{C}(\vec{r},t)=0$. One can therefore decompose the kinetic energy $E$ into two parts, $E=E^{I}+E^{C}$, where
@@
IV.4 Triangular and Honeycomb optical lattice.
We use here the following form:
$V^{tri}_{OL}=\tilde{V}\{\frac{4}{3}-\frac{4}{9}[cos(\vec{b}_{1}\cdot\vec{R})+cos(\vec{b}_{2}\cdot\vec{R})+cos((\vec{b}_{1}-\vec{b}_{2})\cdot\vec{R})]\}$ (16)
$V_{OL}^{hon}(x,y)=-V_{OL}^{tri}(x,y)+2V_{0}$ (17)
We show in Figs.(18) and (19) the equilibrium vortex structure for different values of the potential depth $\tilde{V}$, including the value $\tilde{V}=0$, in the triangular and honeycomb lattice, respectively. In order to facilitate the visualization, as done in Fig.(12), we show the density differences with respect to the configuration without vortex.
Finally, we show in Fig.(20) and (21) the trajectories of a vortex-antivortex pair, obtained by solving the time-dependent equation (2), as they propagate through the lattice before undergoing annihilation after some time. Again, the propagation occurs in the form of jumps across the lower energy barrier sites, and the lower the modulation, the longer it takes for the annihilation to occur and the larger is the displacement of the dipole before annihilation. From these results and those for the square lattice we conclude that the basic features of the motion of a vortex dipole do not depend on the lattice symmetry and that the lifetime of the vortex dipole can be controlled by the intensity of the OL.
IV.5 Visualizing vortices in optical lattices
@@ Acknowledgements.
We thank Gaetano Roati and Alessia Burchianti for useful discussions.

### v3

@@ I Introduction and summary
A somewhat related topics (vortex solitons in BEC stabilized by optical lattices) have been adressed in Ref. malomed, .
@@
A subject that apparently has received very little attention is the behavior of vortex dipoles. Such configurations have been experimentally obtained in BEC in the work of Refs. neely, . However topics such as the interaction of such dipoles and the ensuing dynamics are still unexplored. Our study provide an insight into these issues, and how they are modified by optical lattice.
@@
Vortices in the SS state of matter in dipolar bosons have been addressed gallemi ; casotti ; gallemi1 ; Anc21 in recent years and some relevant phenomena have been uncovered, like a reduced angular momentum associated to the quantum of circulation gallemi . A peculiar behavior has been found Anc21 in the dynamic of a vortex dipole, i.e. a pair of vortices of opposite chirality, in a SS. In a uniform superfluid, when the distance $l$ between the two vortices is larger than the healing length $\xi$, the pair is a stable entity and moves with a constant velocity as a rigid body in a direction perpendicular to the vector $\vec{l}$ joining the two vortices. Contrary to this, in dipolar bosons in a SS state it was found that the two vortices moves by jumps between equilibrium sites in the lattice and approach each other until the two vortices annihilate themselves in a short time. At the basis of this behavior is a basic property of a vortex in a SS: it is energetically favorable to have its core at a discrete set of positions, the locus of minimum density. The natural question is if such properties are specific of a SS state or if they are generic ones when a density modulation is present, whatever is its origin. To answer this question we address in the present paper the study of bosons in an external periodic potential like that produced by light standing waves. We study the ground state and vortices of the standard model of BEC, point-like bosons with a contact interaction, in the case of 87Rb with the Gross-Pitaevskii equation in an external periodic potential $V_{ext}(\vec{r})$. We consider the case of modulation in one dimension (1D) and in two dimensions (2D) for three lattices: square, triangular and honeycomb.
@@
If the two vortices of the dipole are located in the same channel they do not translate but they move one against the other until they annihilate and the excitation energy goes into phonons of the superfluid. If the two vortices are located in adjacent channels they move along these channels and the motion is a composition of a uniform translation and of an oscillation. In all cases the translation velocity of the pair differs from that expected for a uniform superfluid and it can even invert the direction of motion for large amplitude of the optical potential. A combination of translation and oscillation is found also for a single vortex in a trap in presence of a 1D OL. The frequency of this oscillation is the same as in the case of a vortex dipole moving along two neighboring channels. Therefore this oscillatory motion seems to be an intrinsic and novel character of the motion of a vortex moving along a channel. Notice that this oscillatory motion is quite different from that of a massive particle around a minimum energy position because the vortex oscillation is one-side, i.e. the vortex never crosses the position of the minimum energy and it remains on the side from which it started.
@@ II Method
where $V({\bf r})$ and $\rho({\bf r})=|\psi({\bf r})|^{2}$ represent the external potential and the boson number density, respectively. The coupling constant is $g=4\pi a_{s}\hbar^{2}/m$, $m$ being the atomic mass. The number density $\rho$ is normalized such that $\int_{V}\rho({\bf r})\,{\rm d}{\bf r}=N$ where $N$ is the total number of atoms. As model system we take 87Rb atoms and the scattering length $a_{s}$ describing the (repulsive) Rb-Rb interaction is $a_{s}=100.4\,a_{0}$ marte ($a_{0}$ being the Bohr’s radius).
@@
Alternatively, one could estimate the superfluid fraction using Leggett’s formula leg70 :
@@
where $<n(x)>$ is the number density of the ground-state configuration averaged over the transverse y-z directions. The above quantity represents a rigorous upper bound to the superfluid fraction, and it was shown recently perez that under conditions relevant for most ultracold experiments the two definitions (5) and (6) provide surprisingly close values for the superfluid fraction. We indeed find that the values of $f_{s}$ from the two definitions above agree with each other to within $1\%$, in agreement with the findings in Ref. perez, .
@@ III.1 Single vortex properties
The flow field of a linear vortex has a long-range character, $\sim 1/r$, where $r$ is the distance from the position of the vortex core. We have imposed antiperiodic boundary conditions Pi07 in order that the condition of no flow across the boundary of the computational cell is satisfied Anc21 . This is equivalent to sum over the phases of an infinite array of vortex-antivortex, i.e. a vortex of opposite chirality is present in each nearest neighbor cell of the computation cell sadd . Equation (7) can be easily generalized to accommodate a vortex array made of an arbitrary number of vortices and/or antivortices Anc18 . The case of a pair of vortices with opposite circulation (vortex dipole) will be considered in the following. In the case of a vortex dipole we impose however the usual periodic boundary conditions since the flow fields of the two vortices tend to cancel at the cell boundaries.
@@
The equilibrium vortex structure at one of such sites is shown in Fig.3 for the case $\tilde{V}=2$. The streamlines for the velocity field (lines that are tangential to the instantaneous velocity direction) $\vec{v}=\hbar\nabla\phi/m$ ($\phi$ being the phase of the wavefunction) are shown in Fig.4.
Notice the strong deformations with respect to the circular patterns expected from a vortex in a homogeneous system. In order to display these deformations more clearly in Fig.5 we also show the streamlines of the modified phase ${\tilde{\phi}}$ as defined in Eq.(8) further on. ${\tilde{\phi}}$ represents the deviation of the phase from the imprinted phase in Eq.(7).
@@ III.2 Vortex dipole properties
We note that small-amplitude oscillations superimposed to vortex trajectories has been found in a two-component Bose mixture richaud , within the so-called massive point-vortex model, for a single vortex with a rigid circular boundary, where a massless vortex can only precess uniformly. These phenomenology and the one unveiled by our calculations certainly share some common physical properties. The model of Ref. richaud, is based on a Lagrangian formulation for point-vortices, and has been applied so far to homogeneous systems richaud ; richaud1 . It might be of great interest to extend such model to the case of non-homogeneous systems like the one described here.
@@
From the equation for the velocity of a vortex within the GP equation we can understand at a qualitative level some aspects of our results. The flow field of one vortex on the other is in the transverse direction of the channel, i.e. in the region where the density of the superfluid is increasing. This variation of density gives rise to a component of the velocity of the vortex in the longitudinal direction along the bottom of the channel. Following Ref. simula, (see also Ref. sep_joss_rica, ) one may write the vortex wave function $\psi_{v}(\vec{r},t)$ with the core located at position $\vec{r}_{0}$ at time $t$ as
@@
(here $\vec{r}_{0}=(x_{0},y_{0})$ is expressed in complex notation as $z_{0}=x_{0}+iy_{0}$), where the factor $z=x+iy-x_{0}-iy_{0}$ represents the ideal gas vortex wave function and $\tilde{\rho}(\vec{r},t)$ and $\tilde{\phi}(\vec{r},t)$ represent the deviations of the modulus and of the phase from this ideal gas form (see Ref. simula, for additional details). Several effects contribute to the deviation of the vortex wave function from the ideal gas form, i.e. $\tilde{\rho}(\vec{r},t)$ and $\tilde{\phi}(\vec{r},t)$ have the contribution due to the presence of other vortices, to the effect of an external potential and, finally, to the deviation of $\psi_{v}(\vec{r},t)$ from the ideal gas form caused by the inter-atomic interactions. On the basis of the Gross-Pitaevskii equation here adopted the velocity of the vortex is simula
$\vec{v}=\frac{\hbar}{m}(\nabla\tilde{\phi}-\hat{\kappa}\times\nabla log\tilde{\rho})$ (9)
@@
It should be noticed at this point that Eq.(9) is not a self-contained equation for the dynamics of a vortex because the tilded quantities are just a different way to represent the solution of the time-dependent GP equation. As such, Eq.(9) can help to develop qualitative arguments as exposed above, but it cannot predict the dynamics unless approximations are introduced.
@@ III.3 Single vortex in a trap
$V_{box}(x,y)=U_{0}\left[1-\frac{1}{e^{(R-R_{0})/\sigma}+1}\right]$ (10)
@@
We perform the computation with the external potential $V(\vec{r})=V_{OL}(x,y)+V_{box}(x,y)$ and $V_{OL}$ is such that the center of the trap is a maximum of the OL, i.e. the center of the trap is an equilibrium position for the vortex core. We find that the vortex remains immobile if initially it is located at or very near the center of the trap, in the low density channel passing through it. If the initial position of the vortex is off-center along the y direction by a larger amount, however, the vortex is found to migrate towards the edge of the trap and disappears once it approaches the trap edge.
@@ IV.3 Vortex dipole properties
The energy released immediately after the annihilation goes into excitations of the system. In order to gain some insights on the character of the excitations we computed the spectral density of the kinetic energy of the superfluid velocity field, decomposing it into compressible and incompressible parts tsubota ; spectra ; bradley . Briefly, one splits the density-weighted velocity field $\vec{u}(\vec{r},t)\equiv\sqrt{\rho(\vec{r},t)}\vec{v}(\vec{r},t)$ into a compressible (C) and an incompressible (I) part, $\vec{u}(\vec{r},t)=\vec{u}^{I}(\vec{r},t)+\vec{u}^{C}(\vec{r},t)$, such that $\vec{\nabla}\cdot\vec{u}^{I}(\vec{r},t)=0$ and ×$\vec{\nabla}\times u^{C}(\vec{r},t)=0$. One can therefore decompose the kinetic energy $E$ into two parts, $E=E^{I}+E^{C}$, where
@@
We have also considered two other types of 2D lattices, with triangular symmetry and with the honeycomb structure. Similarly with what we have done in the case of the square lattice, we studied the vortex structures in such lattices and the real-time dynamics of a vortex-antivortex pair, eventyally leading to annihilation as the vortices meet. The results are reported in the Supplemental Material SM .
IV.4 Visualizing vortices in optical lattices
@@ Acknowledgements.
We thank Giacomo Roati, Alessia Burchianti and Andrea Richaud for useful discussions.
Supplemental Material
VI Triangular and Honeycomb optical lattice.
We use here the following form:
$V_{OL}^{tri}=\tilde{V}\{\frac{4}{3}-\frac{4}{9}[cos(\vec{b}_{1}\cdot\vec{R})+cos(\vec{b}_{2}\cdot\vec{R})+cos((\vec{b}_{1}-\vec{b}_{2})\cdot\vec{R})]\}$ (17)
$V_{OL}^{hon}(x,y)=-V_{OL}^{tri}(x,y)+2V_{0}$ (18)
We show in Figs.(19) and (20) the equilibrium vortex structure for different values of the potential depth $\tilde{V}$, including the value $\tilde{V}=0$, in the triangular and honeycomb lattice, respectively. In order to facilitate the visualization we show the density differences with respect to the configuration without vortex.
Finally, we show in Fig.(21) and (22) the trajectories of a vortex-antivortex pair, obtained by solving the time-dependent GP equation, as they propagate through the lattice before undergoing annihilation after some time. Again, the propagation occurs in the form of jumps across the lower energy barrier sites, and the lower the modulation, the longer it takes for the annihilation to occur and the larger is the displacement of the dipole before annihilation. From these results and those for the square lattice we conclude that the basic features of the motion of a vortex dipole do not depend on the lattice symmetry and that the lifetime of the vortex dipole can be controlled by the intensity of the OL.